Explain why f(x) = x^2+x+1 is a unit in Z2[x] / (x^3+x+1) and find its inverse.
Answer #1

x txt1 fix) = q*+x+4 € 741x)/(2*2 241) Consider 9x)= x² clearly g(x) E TK213 / 2+x+1) now, fxg6) = 224x+841) = x + 3+ 4+ 224